hamidiso23:

Write the equation of a line that is parallel to y= -5/4x + 7 and that passes through the point (-4, 1).

Vocaloid:

you want a line that is parallel to y = -5/4x + 7. that means your new line has the same slope as this line. the slope of y = -5/4x + 7 is -5/4 now, you have m = -5/4 use point slope form y - y1 = m(x-x1) plug in the point (-4,1) as your x1, y1 solve for y